A bricklayer and an electrician together spend 90 hours working on a new house. If the bricklayer earns $12 per hour, the electrician earns $16 an hour, and the owner pays them a total of $1350 for their work, how many hours does each worker spend on the house?
Answer by ptaylor(1894)

Let x=amount of time spent by the bricklayer
And let y=amount of time spent by the electrician
Now we are told the following
x+y=90----------------------------eq1
and
12x+16y=1350 ($ understood)-----------------------eq2
Multiply eq1 by 12:
12x+12y=1080--------------------eq1a
Subtract eq1a from eq2:
4y=270 divide each side by 4
y=67.5 hours----------------------------amount of time spent by the electrician
Substitute y=67.5 into eq1:
x+67.5=90 add 67.5 to each side
x+67.5-67.5=90-67.5 collect like terms
x=22.5 hours-------------------------------time spent by the bricklayer
CK
67.5+22.5=90
90=90
and
67.5*16+22.5*12=1350
1080+270=1350
1350=1350
Hope this helps---ptaylor